The question is: Use the numbers to create a quadratic equation with the solutions x=-1 and x=1/4.
The numbers are: -5, -4, -3, -2, -1, 1, 2, 3, 4, 5
I have to answer the question in the format:
[number] x^2 + [number] x + [number] = 0

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The question is: Use the numbers to create a quadratic equation with the solutions x=-1 and x=1/4.
------------
That's not a question.
----
create a quadratic equation with the solutions x=-1 and x=1/4.
------------
That's a request.
---
y = (x+1)*(x - 1/4) = x^2 + (3/4)x - (1/4) = 0
===========
The numbers are: -5, -4, -3, -2, -1, 1, 2, 3, 4, 5
IDK what relevance these numbers have.
-------------
I have to answer the question in the format:
[number] x^2 + [number] x + [number] = 0

Answer by ikleyn(52780)

.

With the given roots, a quadratic function has an equation 

    y(x) = a*(x-(-1))*(x-1/4),

where "a" is an arbitrary real number.  Since you want the coefficients from the list, take a = 4 or -4.


You will get then

    y(x) = 4*(x+1)*(x-1/4) = (x+1)*(4x-1) = 4x^2 + 4x - x - 1 = 4x^2 + 3x - 1,

or

    y(x) = -4*(x+1)*(x-1/4) = -(x+1)*(4x-1) = -4x^2 - 4x + x + 1 = -4x^2 - 3x + 1.


These two forms give you two quadratic functions that satisfy the imposed conditions.

Solved, explained, answered and completed.